1)
A favorable cost variance occurs when Oa. actual costs are the same as standard costs Ob. actual costs are more than standard costs Oc. standard costs are more than actual costs Od. standard costs are less than actual costs
2)
The Flapjack Corporation had 8,042 actual direct labor hours at an actual rate of $12.00 per hour. Original production had been budgeted for 1,100 units, but only 999 units were actually produced. Labor standards were 7.9 hours per completed unit at a standard rate of $13.14 per hour. The labor rate variance is O$9,167.88 unfavorable O$9,167.88 favorable $2,153.46 unfavorable O$2,153.46 favorable
3)
If the actual direct labor hours spent producing a commodity differs from the standard hours, the variance is a Oa. time variance Ob. quantity variance Oc. rate variance Od. price varianc
4)
The following data relate to direct labor costs for the current period: Standard costs 36,000 hours at $22.00 35,000 hours at $23.00 Actual costs What is the direct labor time variance? Oa. $23,000 favorable Ob. $35,000 unfavorable c.$36,000 unfavorable Od. $22,000 favorable
5)
The following data is given for the Stringer Company: Budgeted production Actual production Materials: Standard price per ounce Standard ounces per completed unit Actual ounces purchased and used in production Actual price paid for materials Labor: Standard hourly labor rate Standard hours allowed per completed unit Actual labor hours worked Actual total labor costs Overhead Actual and budgeted fixed overhead Standard variable overhead rate Actual variable overhead costs 26,000 units 27,500 units $6.50 8 228,000 $1,504,800 $22 per hour 6.6 183,000 $4,020,000 $1,029,600 $24.50 per standard labor hour $4,520,000 Overhead is applied on standard labor hours The direct materials price variance is Oa. $22,800 unfavorable Ob. Oc. $52,000 favorable $52,000 unfavorable d. $22,800 favorable
Answer #1

1)Answer :C

Explanation : When standard cost is more than actual cost then the variance is said to be favorable.

2)Labor rate variance =( Standard rate- Actual rate ) x Actual hours worked

=(13.14-12)*8042

=9,167.88 favorable

3)Answer :A

Explanation :The direct labor time variance is the difference between the actual direct labor hours and the standard direct labor hours

4)Labor time variance = (Standard hours-Actual hours )* Standard rate

= (36000-35000)*22 per hour = -$22000 favorable

5)Material price variance =( Standard price-Actual price ) x Actual quantity used

=(6. 5-(1504800/228000)) *228000

=22,800 Unfavorable

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW

PLEASE RATE MY ANSWER BY HITTING ??

THANK YOU